OpenStudy (anonymous):

derive (-3x)^(2x) cos(3x) - Sin(2x)

OpenStudy (anonymous):

what do you mean derive

OpenStudy (anonymous):

i ment the derivative of that equation :)

OpenStudy (anonymous):

is it supposed to be = to sin(2x), instead of -?

OpenStudy (anonymous):

you have to derive something from something else

OpenStudy (anonymous):

no that's really -

OpenStudy (anonymous):

erm i don't know. the question doesn't make any sense

OpenStudy (anonymous):

it's like saying derive x + y. you have to derive it from something else

OpenStudy (anonymous):

you see, i'm trying to solve its roots using the newton-rhapson method and i have to get the derivative of \[f(x)=-3x ^{2x}\cos(3x) - \sin(2x)\]

OpenStudy (anonymous):

:( i don't know that method sorry

OpenStudy (anonymous):

that's ok... thanks

OpenStudy (jack1):

\[\large if~y = (-3x)^{2x} \cos(3x) - \sin (2x)\] \[\large then~y' = 9^x (-x)^{2 x} (2 (log(-3 x)+1) cos(3 x)-3 sin(3 x))-2 cos(2 x)\] use chain rule and product rule.

OpenStudy (jack1):

assume you already know base rule: derivative sin = cos, derivative cos = -sin? yeah?

OpenStudy (jack1):

*base rules

OpenStudy (anonymous):

thanks jack1

OpenStudy (anonymous):

jack1? is this answer simplified already?

OpenStudy (jack1):

yeah

OpenStudy (anonymous):

how do you input sec in the sci cal?